OpenStudy (anonymous):

each set of ordered pairs represents a function. Write a rule that represents the function. 1. (0,0) (1,3) (2,6) (3,9) (4,12) 2. (0,-8) (1,-7) (2,-6) (3,-5) (4,-4) 3. (0,-7) (1,-2) (2,3) (3,8) (4,13) 4. (0,8) (1,6) (2,4) (3,2) (4,0)

OpenStudy (anonymous):

f(x) = 3x Ordered pairs are written as (x,y) The way I like to do these is write the y's out as they correspond to x, so write the y for x=0, then for x=1, then x=2 and so on... Find out how to get from one term to the term on its right using addition, subtraction, multiplication, or division If you use + or -, it is an arithmetic sequence (like this one is) An arithmetic sequence can always be written in y=mx+b format The number that you have to add or subtract to get from one term to the next is your m (in this case 3) Then take your x value for each ordered pair and multiply it by m, whatever number you have to add to it to get the y value of the ordered pair is going to be your b In this case, b=0, but b does not always equal zero

OpenStudy (anonymous):

f(x) is the same thing as y, but f(x) usually indicates a function, so you're going to write your answer in f(x)=mx+b format instead of y=mx+b format (even though they mean the same thing)

OpenStudy (anonymous):

agh still makes no sense to me

OpenStudy (anonymous):

I made an x/y chart and tried to explain it better in the attached doc, hope this helps

OpenStudy (anonymous):

This is essentially a "find the pattern question." You can make tables using the ordered pairs and see what the relationship between the x and y values is. Question 1: x | y ---- 0 | 0 1 | 3 2 | 6 3 | 9 4 | 12 Note how this table is essentially just a different way of listing the ordered pairs provided by the question. Taking a quick glance at this table, you should see that any value in the y column is equal to triple the x value in the same row. 3 is triple 1, 6 is triple 2, 9 is triple 3, etc. Knowing this, write an equation (a "rule") to represent the relationship. In this case, the y value is triple the x value, so the rule is that y = 3 * x. This approach should allow you to solve the rest of the problems. You may even be able to imagine the table in your head and come up with a rule without writing anything down. Please let us know if you need more advice. Also, let us know what you get for answers to the questions so we can make sure you understand.